OpenStudy (hoothoote):

Help! I'll fan and medal! f(x) = 3x + 1; g(x) = 5x - 1 Find f/g

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (hoothoote):

Answers: Find f/g. (f/g)(x) = Quantity five x minus one divided by three x plus one.; domain {x|x ≠ - One over three. } (f/g)(x) = Quantity three x plus one divided by five x minus one; domain {x|x ≠ - One over three. } (f/g)(x) = Quantity five x minus one divided by three x plus one.; domain {x|x ≠ One over five. } (f/g)(x) = Quantity three x plus one divided by five x minus one.; domain {x|x ≠ One over five. }

OpenStudy (hoothoote):

@rebeccaxhawaii

Vocaloid (vocaloid):

look at what rebecca wrote and see which answer choice matches.

Vocaloid (vocaloid):

don't forget to solve for excluded x values

OpenStudy (hoothoote):

Either B or D but I'm not sure how to solve the second part

Vocaloid (vocaloid):

set the denominator equal to 0 5x - 1 = 0 now solve for x.

OpenStudy (hoothoote):

Is the second part 1/5? @Vocaloid

Vocaloid (vocaloid):

good

Vocaloid (vocaloid):

so your domain is domain {x|x ≠ One over five

OpenStudy (hoothoote):

Thank you!